Storico Co. just paid a dividend of $1.65 per share. The company will increase its dividend by 24 percent next year and then reduce its dividend growth rate by 6 percentage points per year until it reaches the industry average of 6 percent dividend growth, after which the company will keep a constant growth rate forever. If the stock price is $46.84, what required return must investors be demanding on the company's stock? (Hint: Set up the valuation formula with all the relevant cash flows, and use trial and error to find the unknown rate of return.) (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

D0 = $1.65 per share

Dividend will grow at (g1) = 24%

Thereafter the dividend growth rate will start declining by 6% per year until it reaches 6% per year which is industrial average

D1= D0(1+g1) = $1.65*(1+24%) = $2.046

g2 = 18% (24%-6%)

D2= D1(1+g2) = $2.046(1+18%) = $ 2.414

g3 = 12%

D3 = D2(1+g3) = $2.414(1+12%) = $2.704

g4 = 6% (now growth will be constant as it has reached the industrial average)

D4 = D3(1+g4) = $2.704(1+6%) = $ 2.866

Now, calculating required return(ke) by using IRR technique in dividend discount model;

D1 D Po = (1 +he)i + (1+ke)2 + (1+ke)3 + (1+ke)3 * (he – ga

2.046 2.414 2.704 2.866 PO = (1 + ke)1 + (1+ke)2 + (1+ke)3 + (ke -0.06) * (1 + ke)3

Taking ke = 10%

Po= 2.046 2.414 2.704 2.866 0 = (1 +0.10)1 *(1 +0.10)2*(1 +0.10)3 * (0.10 – 0.06) * (1 +0.10)3

P0 = $59.7183

taking r = 11%

Po = 1 2.046 2.414 2.704 0(1+0.11)1 (1+0.11)2 *(1 +0.11)3 2.866 (0.11 -0.06) *(1 +0.11)3

P0 = $ 47.6915

Now, calculating ke

PriceLR – Current Price Ke= Lower Rate+ – **(Higher Rate-Lower Rate) PriceLR - Pricehr

Ke=10+ 59.7183 - 46.84 - * (11 – 10) 59.7183 - 47.6915*

12.8783 Ke = 10+ 12.0268 * (11 - 10)

ke = 11.10%

So, required return must investors be demanding on the company's stock is 11.10%